### v1

Effect of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentialsthanks: We would like to thank Sandra Black, Rich Murphy, Steve Trejo, Gerald Oettinger, Dan Hamermesh, Kitt Carpenter, Dario Sansone, Chris Jepsen, Luis Faundez, and all the seminar participants of the University of Texas at Austin. No financial support was given to support this research. Correspondence: d.vamossy@pitt.edu.
Scott Delhommer University of Texas at Austin; sdelhommer@utexas.edu    Domonkos F. Vamossy University of Pittsburgh, d.vamossy@pitt.edu
(April 4, 2024)
This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on local anti-discrimination laws. Using variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws significantly reduce gaps in labor force participation rate, employment, and the wage gap for gay men relative to straight men. These laws also significantly reduce the labor force participation rate, employment, and wage premium for lesbian women relative to straight women. One explanation for the reduced labor supply and wage premium is that lesbian couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
JEL Codes: J31, J71, J78, K31
1 Introduction
On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals. Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT. Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020. As a result, laws extending protected class status based on sexual orientation varied between states and localities. The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act. This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner. There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics. The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers.
We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts222(Badgett (1995); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al. (2003); Carpenter (2005); Black et al. (2007); Antecol et al. (2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)). The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
There have been previous studies examining how these state and local laws affect the pay gap, but these studies have been limited in their ability to identify causal effects (Klawitter and Flatt (1998); Gates (2009)). Klawitter and Flatt (1998) compared people in same-sex partnerships to those in different-sex partnerships in the 1990 Census and found that same-sex couples have no significant difference in wages in places with anti-discrimination laws. Gates (2009) uses the 2000 Census and conducts a similar analysis. Gates (2009) finds that gay/bisexual men in places with anti-discrimination laws have a 3% wage premium over gay/bisexual men in places without these laws, and lesbian/bisexual women have a 2% wage premium over lesbian/bisexual women in places without these laws [^3] Tilcsik (2011) found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws. These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework. [^4] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008). Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.555 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences. Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design. Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape. Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context. Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2004 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S. census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples. The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years. This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples. Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances. Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment. To bridge this gap, our research employs the methodology proposed by Borusyak et al. (2021), comparing it with the conventional triple difference-in-differences estimator, finding both approaches yield similar results.
This paper is the first to examine both local and state-level sexual orientation anti-discrimination laws in a quasi-experimental design. This paper is also the first to analyze the effects of any local anti-discrimination laws on sex, race, or sexual orientation in a modern difference-in-difference framework [^6] Using state variation in anti-discrimination laws has also been analyzed to understand racial- and sex-based discrimination (Neumark and Stock (2006); (Donohue III and Heckman, 1991);((Margo, 1995); Goldin and Margo (1992)). However, these studies on sex and race anti-discrimination laws also omit local laws, potentially skewing their estimates..
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws significantly reduce the gap in labor force participation and employment of gay men by 2.1 p.p. and 1.5 p.p., respectively. The laws also significantly increase hourly wages for gay men by 6.2%. The results differ for women, with the laws significantly reducing their labor force participation and employment premium over straight women by 1.7 p.p., and 2.3 p.p., respectively. Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons. [^7] In Appendix A, we replicate the findings of past research that finds gay/bisexual men have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings. We also replicate findings that lesbian/bisexual women have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
We explore theories for the differing effects of anti-discrimination laws on gay men and lesbian women in the Discussion section, using the Becker (1981) model of household specialization. In the traditional Becker (1981) model of household specialization, men typically specialize in market production, and women typically specialize in household production, in part due to differences in biology where women birth and care for children, resulting in a one-earner household. These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in same-sex partnerships following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that the difference in hours worked between partners within a lesbian household goes up relative to gay households following anti-discrimination laws, suggesting that lesbian households could become more specialized following anti-discrimination laws with one woman working more hours and the other woman working fewer to focus on household production. We also show that lesbian households have significantly more children than gay households after the passage of an anti-discrimination law. More children could induce lesbian households to further specialize in the intra-household division of labor, adopting a more traditional household model to help care for additional children.
Finally, our last contribution is the use of polling data on the support for same-sex marriage to examine the relationship between anti-discrimination laws and state sentiment toward LGB workers. One may expect that sentiment toward LGB workers would increase right before the passage of the laws, creating a selection issue. Alternatively, the anti-discrimination laws may normalize being a sexual minority, and improve sentiments toward LGB workers. We collected every poll on same-sex marriage for each state and year from the Pew Research Center. We show that the passage of state-wide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage. The states that passed anti-discrimination laws had parallel pre-trends in support of same-sex marriage before the passage of the laws. Afterward, those states had a significant and persistent increase in their support for same-sex marriage. The observed increase in support following the enactment of the law, rather than before, implies that policy changes can indeed influence public opinion. This notion aligns with the findings of Sansone (2019), suggesting that legislative actions can lead to shifts in societal attitudes. This is in contrast to the thermostatic model proposed by Wlezien (1995), which posits that public sentiment drives policy changes rather than the other way around. Increased favorability toward LGB people may act in conjunction with greater job security to improve labor market outcomes for LGB Americans.
2 Data
A common issue in analyzing the pay gap/premium between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment. We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition. Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves. [^8] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009).. The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that straight roommates would misclassify as unmarried partners.
The comparisons that we make are between people in same-sex relationships, defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex relationships. This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers. It compares those who are in a same-sex relationship to those who are in a different-sex relationship. Notably, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual people in different-sex relationships as members of the control group. For example, if a person is misclassified as straight instead of bisexual, and bisexual workers have lower pay than their straight counterparts, then the misclassification would bias our pay gap/premium results toward zero.
We use the 2005-2019 yearly ACS from (Ruggles et al., 2019). We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked999“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked.. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. This dataset is the first complete city-level dataset on sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together. We make the assumption in our analysis that the impact of these laws is the same across states and localities, estimating an average effect.
We report the state, city, and county laws and the year they were enacted in Table 1. In Figure 1, we show which counties had sexual orientation anti-discrimination laws in 2005 and 2019, respectively. In our analysis, laws are coded based on their initial application to private employment and housing, focusing on broad protections rather than sector-specific ones. For instance, while Delaware extended protections to state employees in 1999, our coding reflects 2009, the year broader protections were enacted. This approach could overlook the incremental impact of earlier, narrower protections on societal attitudes. The same principle applies to local laws, where initial protections for public employees in some cities might precede more comprehensive legislation, potentially influencing early shifts in public sentiment towards sexual minorities. The enforcement mechanisms of anti-discrimination laws, such as fines, mandatory training, and public rebukes, serve as significant deterrents, encouraging the adoption of more inclusive employment policies. The rise in successful legal challenges against companies concerning discrimination highlights the real legal and financial consequences they encounter, reinforcing the laws’ effectiveness in promoting workplace equality (for Human Resource Management (2023)).
The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^10] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. are noticeably absent from extending protection to LGB workers. One potential concern in this analysis is the endogenous adoption of laws. The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws. The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias. It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes. We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018). The inclusion of county-year fixed effects allows us not to include county controls hence eliminating the need to control for unavailable local levels of support for same-sex marriage. We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), particularly suited for analyzing staggered policy adoptions like ours. This method accurately estimates the effects of anti-discrimination laws, considering their varied implementation times. For additional validation, we include results from a simpler difference-in-differences estimator in the appendix, providing robustness to our findings.
In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships. Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston. The LGB population varies from the straight population in many dimensions including geography.
We present descriptive statistics showing the differences in the LGB population compared to the straight population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
3 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time and between same-sex and different-sex couples lends itself to a saturated triple difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county. Formally, we estimate this equation on labor supply:
$\begin{split}Y_{ict}=\alpha+\beta_{0}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Female})}+\beta_{1}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Male})}\\ +(I_{c}\times I_{t}\times I_{\textrm{Sex=Male}}))\gamma_{sct}+(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}}))\theta_{S,s,c}\\ +(SSP_{ict}\times I_{t}I_{\textrm{Sex=Male}}))\theta_{S,s,t}+X^{\prime}_{ict}\delta+\epsilon_{ict}\end{split}$ (1)
In our regression, equation (1) defines the labor market outcomes $Y_{ict}$ for individual $i$, in county $c$, at time $t$ with sex $s$ and same-sex partnership status $S$. These include the labor force participation, employment status, or the real value of wages, in 1999 dollars. The binary variable $SSP_{ict}$ indicates if an individual is in a same-sex partnership. The model explores the interaction between being in a same-sex partnership and the presence of county-level Employment Non-Discrimination Acts $(SSP_{ict}\times ENDA_{ct})$. [^11] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. (2020), there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The fixed effects for county-year-sex interactions $(\gamma_{ct})$, control for unobserved heterogeneity. Interactions between the same-sex partnership status, sex and county $(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}})$ and year fixed effects $(SSP_{ict}\times I_{t}\times I_{\textrm{Sex=Male}})$ capture varying impacts over time and across different regions by sex. $X^{\prime}_{ict}$ is a vector of person-specific covariates including race, children, age, and education level - each represented up to a quadratic term to capture potential nonlinear effects. The error term $\epsilon_{ict}$ is clustered at the county level to account for within-group correlation. [^12] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties. This specification aims to isolate the effects of anti-discrimination laws on the labor market outcomes for individuals in same-sex partnerships by sex.
Our sample is limited to people aged 25 to 65, and we estimate this model jointly by sex using the ACS person weights. When examining wages, the $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum.
In equation (1), the differential effect of being a cohabiting same-sex couple by sex in the United States over time is captured by $\theta_{S,s,t}$. The county fixed effects for cohabiting same-sex couples by sex $\theta_{S,s,c}$ capture how the wage differential varies across counties by sex. The parameters of interest is $\beta_{0}$ and $\beta_{1}$, which capture how the differential between homosexuals and heterosexuals in counties with ENDAs changed after the law’s passage by sex. Hence, $\beta_{1}$, will give the effect of ENDAs on the labor supply and pay gap/premium for males, and $\beta_{0}$ for females. [^13] We recognize the heterogeneous impact of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
One concern is that the distribution of LGB workers is not random, and states/counties that pass anti-discrimination laws have a higher concentration of LGB workers. However, using county fixed effects will make comparisons within a given county, so San Francisco will be compared to San Francisco. A potential concern would be if there is sorting of LGB workers to areas with anti-discrimination laws protecting them, changing the composition of the area. We show that there is little evidence to support this in the robustness check section.
To interpret $\beta_{0}$ and $\beta_{1}$ as being the causal effect of enacting employment protections for sexual orientation, it must be the case that there are no other factors related to changes in the gay/lesbian-heterosexual wage differential in states that do and do not have LGBT employment protections. To accurately estimate $\beta_{0}$ and $\beta_{1}$ while considering the staggered implementation of ENDAs, we employ the methodology from (Borusyak et al., 2021). This approach uses observations only for units and periods that are not yet treated to infer the never-treated potential outcome for each treated unit.
Our identifying assumption in this framework is that, in the absence of anti-discrimination laws, the trends in labor supply and wage disparities in counties with such laws would have mirrored those without them. This difference-in-differences framework hinges on parallel trends across all groups and periods, assuming no prior anticipation of the laws. While directly testing this parallel trend assumption is challenging, examining pre-treatment trends can provide some indication of its validity. We examine the outcomes relative to the timing of the laws by estimating an event study model, where we estimate coefficients for each period relative to the treatment time. This approach allows us to assess the impact of the laws over time and ensure a thorough understanding of their effects before and after implementation.
We plot the effect of the law on the labor supply and pay gap/premium in each year relative to the start date. The effect of the law in the year before the law is enacted is normalized to zero. We report the coefficients estimated via Borusyak et al. (2021) in the main text, and standard TWFE in the Appendix.
This paper’s main contribution is applying a staggered policy design approach, providing an understanding of the average effects of anti-discrimination laws. Our approach highlights the importance of considering local contexts when assessing the efficacy of legislative measures.
4 Primary Findings
4.1 Extensive & Intensive Margins, Wages
In Table 4, we present the regression on the extensive & intensive margins of labor supply with $\beta_{1}$ giving the effect of anti-discrimination laws on gay men’s labor supply and $\beta_{0}$ giving the effect for lesbian women. Columns (1-2) show that anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men. The labor force participation rate gap is reduced by 2.1 p.p., and the employment gap is reduced by 1.5 p.p. Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.7 p.p. and 2.3 p.p., respectively. Columns (3-6) present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. We find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is insignificantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced when measured in weeks worked. We later show that the positive impact for lesbian women on weekly hours and the negative effect on weeks worked for gay men are gone once controlled for occupational fixed effects.
In Table 5, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a $\ln(Y_{ict}+1)$ transformation, to handle zero values. [^14] We see small differences in effect sizes in a log-only specification on the employed-only population. We run the regressions jointly by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table 4, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers.
Gay men experience a notable increase in hourly wages of 6.2%, which is statistically significant at the 95% confidence level. Additionally, there is a substantial increase of 22.2% concerning their annual earnings, indicating significant progress towards income equality. On the other hand, lesbian women see an insignificant change in their annual earnings premium and an insignificant reduction in their hourly wage premium. The pay gap/premium is reduced quite significantly once the sample is limited to employed workers in columns (3) and (4), which is unsurprising. The effects also turn insignificant, which shows that they come from increased labor force participation and employment rates.
4.2 Occupation Fixed Effects
LGB workers tend to sort to different occupations and industries based on sex and sexual orientation (Black et al., 2007). They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts. It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
While controlling for occupation will bias the wage gap, it can still be instructive for seeing the effect of anti-discrimination laws on the wage and labor supply gap. We run these regressions with occupation fixed effects using four-digit occupation codes in the ACS to see how the pay gap/premium changes after anti-discrimination laws are passed. Using occupation-fixed effects will examine the effect of anti-discrimination laws within an occupation. One potential mechanism for anti-discrimination laws is that gay workers sort into more in-demand occupations, lowering their relative labor supply and wage gap, while lesbian workers do the opposite. However, we show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within an occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table 6 and 7.
In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry. However, qualitatively the results are similar to the main results in Table 4 and Table 5. Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
4.3 Impacts Over Time
We use Equation (1) with the inclusion of occupation-fixed effects and conduct event study regressions via Borusyak et al. (2021). This approach allows us to analyze how the impact estimates evolve. First, we present the event study regressions for male and female extensive labor supply outcomes in Figure 3. Panel (a) illustrates the effects on labor force participation rates, whereas Panel (b) shows the effects on employment levels. The plots reveal an initial jump in both employment and labor force participation rates among gay men, a trend that appears to wane over subsequent years. This pattern hints at an early beneficial impact that seems to moderate with time. Conversely, the trajectory for women follows a different pattern, indicating initial adverse effects on employment and labor force participation rates. The pre-trends look fairly parallel, which, while not directly confirming the parallel trends assumption for the post-period, lends it greater plausibility. These visualizations indicate a temporary increase in labor force participation and employment rates for gay men following the enactment of ENDAs, contrasting with a negative impact on the same metrics for lesbian women.
We present the event study regressions for male and female pay penalty/premium outcomes in Figure 4. The figures suggest a delayed response in wage parity gains for gay employees compared to their heterosexual counterparts, with a notable but transient increase in relative gay wages observed a year post-ENDA implementation. This delayed effect may come from implementation lags and the labor market outcomes captured in the ACS, which reflect earnings over a 23-month reference period. Conversely, there is a marked decline in relative wages for lesbian women following the enactment of the law, becoming significant one year after its introduction.
5 Robustness Checks
5.1 Endogenous Adoption of Anti-Discrimination Laws
One concern in this analysis is the endogenous adoption of anti-discrimination laws. Anti-discrimination laws are not randomly distributed. Locally, anti-discrimination laws are concentrated in larger cities, and state laws are to be concentrated in more liberal states that presumably are more accepting of and more favorable to LGB workers. The main concern is that there is an unobservable factor like general sentiment toward LGB workers that affects both the passage of laws as well as the labor market outcomes for LGB workers. In our main regressions, we alleviate this concern by using county-year-sex fixed effects.
Nonetheless, we investigate the endogenous adoption of anti-discrimination laws at the state level. Specifically, we use state-level polling information on support for same-sex marriage as a proxy for general sentiment toward LGB workers. It is possible that controlling for state-level polling is not the best way to capture sentiment toward LGB workers since it is possible to discriminate against people based on sexual orientation and still support their right to marry. However, it seems plausible that the changes in state-level support for same-sex marriage are highly correlated with changes in sentiment toward LGB workers such that it will suffice for a suitable proxy. To better get at the question of endogenous adoption of laws, we create an event-study plot showing how state laws change support for same-sex marriages. Specifically, we estimate this equation:
$\text{Support}_{st}=\sum_{j=-5,j\neq-1}^{j=5}\psi_{j}\mathbbm{1}(\text{Laws}_{st}=j)+\phi_{s}+\delta_{t}+\epsilon_{ts}$ (2)
[…]
The event study supports the idea that anti-discrimination laws at the state level significantly increase the support for same-sex marriage at the state level. The trends before the legislation exhibit similar trends, with a noticeable jump in the fraction of individuals favoring same-sex marriage. This pattern suggests that states, irrespective of their legal stance on anti-discrimination, shared similar trajectories in support for same-sex marriage before the enactment of such laws, and only began to diverge following the implementation of the legislation. It might be argued that the passage of these laws was a consequence of shifting attitudes towards same-sex marriage at the time of enactment. Still, if the adoption of these laws were endogenous, one would expect to see a steady climb in support for same-sex marriage leading up to the law’s passage, with the legislation itself having minimal impact on this trend. However, the distinct and significant jump in support concurrent with the introduction of anti-discrimination laws lends more weight to the argument that it is the legislative changes that are driving the increase in support for same-sex marriage and, likely, increasing the overall sentiment towards LGB individuals, rather than the other way around.
The jump seen in state-level sentiment toward LGB workers following state anti-discrimination law passages could suggest that there is a corresponding jump in sentiment at the local level that is unobservable and potentially acting as a mechanism, affecting the labor supply and pay for LGB workers.
5.2 Sorting and Increased Reporting
One finding from Klawitter and Flatt (1998) suggests that LGB workers sort to areas with anti-discrimination laws. If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption. It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
Another potential concern would be increased reporting for same-sex partnerships. Same-sex partnerships, while less stigmatized in 2005-2019 than in previous years, were still heavily stigmatized. Anti-discrimination laws could signal to LGB people that their community is more accepting, and people may be more likely to declare that they are in a same-sex partnership. This change in reporting could affect the composition of our sample, which in turn could violate the identifying assumption.
To address this concern, we run the regression below to see if there is an increase in same-sex partnerships in a given county after an anti-discrimination law is passed:
$\text{SSP}_{ict}=\beta_{0}\text{Laws}_{ct}+\beta_{1}X_{i}+\beta_{2}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (3)
In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships. We present these results in Table 8.
We find no significant effect of anti-discrimination laws on the number of same-sex partnerships at the county level, suggesting there is minimal sorting or change in reporting following the passage of anti-discrimination laws.
6 Discussion
6.1 Differences in Response by Sex
The results paint an interesting picture of the effect of anti-discrimination laws on the workforce, in that they have differing effects on lesbian and gay workers. Lesbian workers see a reduction in the labor supply and pay premium while gay workers see a reduction in their labor supply and pay gap. The anti-discrimination laws appear to push the labor market outcome to a greater level of equality across sexual orientation, where lesbian women move toward the labor market outcomes of straight women and gay men move toward the labor market outcomes of straight men.
One potential explanation for this convergence of labor market outcomes is that employers stop differentiating between workers based on sexual orientation after anti-discrimination laws are passed and that results in greater equality. However, much of the labor market differences arise from differences in human capital accumulation and decisions that come before many of the anti-discrimination laws such as occupation/industry choice and desired number of children (Black et al., 2007).
Another explanation for different responses to anti-discrimination laws for same-sex couples lies in the canonical Becker (1981) theory of the family and division of labor, which is typically used to explain differences in LGB labor market outcomes from their straight counterparts (Black et al., 2007). In a male-female household with a “traditional” division of labor, the man specializes in market production, selling his labor for money to buy market goods, and the woman specializes in household production, using her labor to rear children, cook, and take care of the home. However, LGB people are less likely to end up in a traditional Beckerian household. This expectation of future household composition will induce gay men to be less specialized in market production and lesbian women to be more specialized in market production than their straight counterparts. This theory of household specialization is typically used to explain why gay men make less than straight men and lesbian women make more than straight women (Black et al., 2007).
This household model can also be used to think about responses to anti-discrimination laws. One potential explanation for lesbian women becoming more likely to leave the workforce and earn less is that a lesbian household may be more likely to shift to a more traditional Beckerian household where one parent specializes in household production and the other shifts to a more market orientation. If anti-discrimination laws decreased the risk of the primary earner getting fired for their sexual orientation, it could induce the secondary earner to switch to more part-time work or exit the workforce to specialize in household production or focus on child-rearing. Alternatively, both women in a female same-sex partnership could switch to reducing hours to spend more time in household production as a result of increased job security.
Households consisting of lesbian women typically have more children than households consisting of gay men. Given lesbian women have more children than gay men, it would be unsurprising to see lesbian households switch more to a specialization of their household labor than gay men would, which would help explain why anti-discrimination laws have different effects on lesbian women and gay men.
To test the different responses by men and women in same-sex partnerships to anti-discrimination laws and the implications of Becker’s household model, we implement an alternative difference-in-difference model. Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level. Do women in same-sex partnerships change their household division of labor relative to men in same-sex partnerships following anti-discrimination laws? To answer this question, we collapse the data to the household level and examine households instead of individuals.
Specifically, we estimate this model:
$y_{jct}=\alpha_{0}Law_{ct}+\alpha_{1}FemSSP_{j}+\alpha_{2}Law_{ct}*FemSSP_{j}+\mu_{c}+\delta_{t}+\epsilon_{jct}$
The notation is the same as before with two main differences. Our interaction term uses an indicator for if household $j$ is a female same-sex partnership, denoted by $FemSSP_{j}$. $y_{jct}$ is a variable for outcomes at the household level. $\alpha_{2}$ is our coefficient of interest and will show how lesbian households differ in their response to anti-discrimination laws compared to gay households. The first outcome examined is an indicator for if the household is a one-earner family. We also examine the difference in absolute value between the two partners in terms of weekly hours worked. These labor supply measures inform how intra-household labor is divided between partners along the extensive and intensive margin of labor supply. Finally, we examine how the anti-discrimination laws may affect child rearing by examining if the households have any children and the number of children they have.
We present the results in Table 9. Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin. There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for gay and lesbian households. An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership. Lesbian households do see a significant increase in the likelihood of having any child and the number of children they have relative to gay households.
This evidence tied with Becker’s theoretical work suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws. Lesbian households begin to have more children than gay households. It may appear in the main results that anti-discrimination laws hurt lesbian women because their labor supply and pay premium over straight women shrinks. However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security. The corresponding changes in the labor market could represent an increase in welfare for lesbian households if they value their new children more than their lost hours worked and subsequently lost wages. Consistent with the theory and empirical findings, lesbian households could become more secure with one woman being the primary earner following anti-discrimination laws.
6.2 Mechanisms
The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for gay/bisexual men as well as the labor market premiums for lesbian/bisexual women. It is possible that these laws are binding and eliminate significant discrimination in the labor market. However, it is relatively easy for an employer to create a reason to fire an employee rather than firing someone specifically for being gay. It is also difficult to accurately determine someone’s sexual orientation by looking at them, unlike race and sex. Although it may be easier to determine someone’s sexual orientation conditional on them being in a same-sex relationship. Sexual orientation anti-discrimination laws may not be binding in a meaningful way that eliminates discrimination.
One mechanism that the laws could be affecting labor market outcomes is through increased positive sentiment toward LGB workers. We show in Figure 5 that following state anti-discrimination laws there is a significant and persistent increase in the percentage of people in that state that support same-sex marriage. Our regression results already condition on the change in the county-level support for same-sex marriage via fixed effects. In Appendix A, we show that local laws matter significantly, so it seems plausible that local law changes could significantly change local sentiment and impact the labor supply and pay of LGB workers. Changing public sentiment in conjunction with greater protection for LGB workers seems more plausible as the mechanism than simply greater protection in the workplace.
6.3 Threats to External Validity
One limitation of this research is that we are unable to view all LGB workers. The analysis is specifically conducted on individuals in same-sex partnerships and different-sex partnerships. The data does not allow for the identification of single LGB workers, whose outcomes may be significantly different than LGB workers in partnerships. Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership. Bisexual people are a larger group than lesbian or gay people, and this specification could erroneously assign bisexual people to be “straight” through being in a different-sex partnership (Gates, 2011). If bisexual workers appear more similar to their lesbian and gay counterparts then the results of quantifying the labor supply and pay gap/premium would likely be biased toward zero and the change in those labor market gaps/premiums would also be biased toward zero.
Another potential concern in using partnerships to infer sexual orientation is that people in a same-sex partnership may be the group that is most at risk of discrimination. It seems likely that single LGB workers can more plausibly stay in the closet with their co-workers compared to their counterparts in a same-sex partnership. This analysis may be capturing the effect of anti-discrimination laws on the group that is most likely to be affected. These results may have a larger effect size than what one would find in examining the generalized LGB population.
These biases are for the broader LGB population and external validity. Notably, these potential biases do not affect internal validity. Translating the effect of anti-discrimination laws for people in same-sex partnerships to all other LGB workers is not immediately obvious. It is dependent on several factors that are unknowable in this analysis such as the wages and labor supply for single LGB workers and bisexual workers in different-sex partnerships and how discrimination affects LGB people for those in same-sex partnerships.
7 Conclusion
This analysis is the first quasi-experimental research examining how both local and state anti-discrimination laws on sexual orientation affect the labor supply and pay gap/premium between LGB and straight workers. We construct a novel panel dataset on local anti-discrimination laws to capture the granular nature of anti-discrimination laws. We find that anti-discrimination laws appear to decrease sexual orientation inequality in the labor market. Specifically, we find that anti-discrimination laws significantly reduce the gay labor force participation gap by 2.1 p.p., the employment gap by 1.5 p.p., and increase hourly wages by 6.2%. Anti-discrimination laws also reduce the lesbian labor force participation premium by 1.7 p.p., and the employment premium by 2.3 p.p. We also show that one potential mechanism that anti-discrimination laws work through is increasing positive sentiments toward LGB Americans as measured by the support of same-sex marriage. Finally, we explain the differential response to these laws between gay and lesbian households through Becker’s household specialization model and support the explanation with empirical evidence.
Before 2020, it was legal in some states to fire someone solely based on their sexual orientation, and states have been less inclined to pass anti-discrimination laws recently. Arkansas and Tennessee even passed anti-anti-discrimination laws preventing cities from enacting anti-discrimination laws. From 2009 to 2019, more states have barred cities from protecting their residents than the number of states extending protection. Our research gives a comprehensive look at a policy that Congress is considering in “The Equality Act” to extend federal protection to sexual orientation and gender identity. Our research gives the most comprehensive look at policy and can meaningfully inform the policy discussion around anti-discrimination laws. Overall, this research broadly informs the policy decision of extending employment protections to the LGB community by the Supreme Court of the United States.
Appendix A TWFE Estimates
In the original version of our paper, we employed a traditional difference-in-differences (DiD) setup, incorporating a range of controls to assess the impact of anti-discrimination legislation on labor market outcomes for LGB workers. This conventional approach served as the foundational analysis for our initial findings. To ensure transparency and thoroughness in our research, we are including these initial estimates in this section. Furthermore, in response to recent advancements in DiD methodologies, we use Borusyak et al. [2021] to estimate the impact of anti-discrimination laws. Our analysis reveals that the application of these modern DiD techniques does not substantially alter the point estimates we initially obtained. This consistency across methodological frameworks underscores the robustness of our findings, reinforcing the reliability of our conclusions regarding the effects of anti-discrimination laws on LGB labor market outcomes.
In the TWFE setup, we control for sentiment toward LGB workers in a given state at a given year by including polling information on support for same-sex marriages. Unfortunately, we cannot observe polling at a local level, as such the polling measure is an incomplete proxy for sentiment toward LGB workers. We collected every poll that Pew Research Center published from 2005-2016 to pull each poll regarding support for same-sex marriage. The 28 polls asking about same-sex marriage were aggregated to get the percentage of people that supported same-sex marriage by state by year. The polling information contains data on every state for every year, except for Alaska and Hawaii, which are missing polling information for 2005-2008.
A.1 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, LGB workers and heterosexual workers. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county.
Formally, we estimate this equation on labor supply:
$LS_{ict}=\alpha_{0}Law_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}Law_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
Where $LS_{ict}$ is the labor force participation or employment status for person $i$ in county $c$ in year $t$. $Law_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership. $X_{i}$ is a vector of person-specific covariates including race, children, along with age and education level - each represented up to a quadratic term to capture potential nonlinear effects. $\gamma_{st}$ is a vector for state $s$ covariates on the legality of same-sex marriage in year $t$ as well as polling for the support of same-sex marriage. It includes the percentage of people that “strongly support,” “support,” “oppose,” and “strongly oppose” same-sex marriage in a given state in a given year. Finally, $\mu_{c}$ gives county fixed effects, $\delta_{t}$ gives year fixed effects, and $\epsilon_{itcs}$ is the error term [^15] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties.. Our sample is limited to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. We cluster the standard errors at the county level, as that is the level of the treatment in this specification. [^16] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. [2020], there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The coefficients of interest are $\alpha_{2}$, which will give the effect of the law on the labor supply gap/premium, and $\alpha_{1}$, which gives the labor supply gap/premium in the absence of anti-discrimination law. Our specification assumes that the effect of the laws is the same across states and localities, and we will be estimating an average effect. We estimate a similar equation for wages:
$\ln(Y_{ict}+1)=\beta_{0}\text{Law}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{Law}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
Where $Y_{ict}$ is the real value, in 1999 dollars, of person $i$’s wages in county $c$ year $t$. The rest of the notation is the same as above. The coefficients of interest are $\beta_{2}$, which will give the effect of the law on the pay gap/premium, and $\beta_{1}$, which gives the pay gap/premium. [^17] The $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum, capturing nuances missed by focusing solely on employed workers. We report the coefficients estimated via Borusyak et al. [2021] in the main text, and standard TWFE in the Appendix.
A.2 TWFE Results
In Table A1, we present the regression on the extensive margin of labor supply with Panel A giving the effect of anti-discrimination laws on gay men’s labor supply and Panel B giving the effect for lesbian women. We present the full specification in columns (1) and (2) and include a specification without accounting for local anti-discrimination laws in columns (3) and (4) to show how failing to account for local laws would lead to incorrect inference.
Consistent with the previous literature, we find that lesbian women have a labor supply premium and gay men have a labor supply penalty on the extensive margin. Lesbian women are 7.3% more likely to be employed and gay men are 7.9% less likely to be employed than their straight peers. Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men. The labor force participation rate gap is reduced by 1.3 p.p. (18%), and the employment gap is reduced by 1.4 p.p. (18%). Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.3 p.p. (16%) and 1 p.p. (14%), respectively. As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws only using state laws will lead one to erroneously conclude that these laws have no significant effect on the gay labor force and employment gap. However, the effect on the extensive margin of labor supply for lesbian women looks similar when estimating the effect of just state anti-discrimination laws.
In Table A2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. Similar to Table 4, we include the results when only estimating the effects of the state laws in columns (3) and (4) to show the contribution of including local laws in estimating the effect of anti-discrimination laws.
We replicate past findings where gay men supply significantly less labor and lesbian women supply significantly more labor on the intensive margin. As with Table A1, we find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced as well. The effect of anti-discrimination laws on weekly hours worked becomes insignificant for gay men and lesbian women when only estimating off of state-wide anti-discrimination laws.
In Table A3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a log-transformation, to handle zero values so the changes can be interpreted in percentage terms. We run the regressions separately by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table A1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers. We include results by only using state anti-discrimination laws, excluding local laws. We present those results in columns (5) - (8).
Gay men see a significant decline in the hourly wage gap by 2.8 p.p. (11%) at the 90% significance level and see an insignificant decline in the annual earnings gap by 5.8 p.p. On the other hand, lesbian women see a significant change in their annual earnings premium by 13 p.p. (16%) and an insignificant reduction in their hourly wage premium. Columns (3) and (4) replicate the past literature where employed gay men make about 8% to 11% less than their straight counterparts and employed lesbian women make about 5% to 15% more than their straight counterparts.
The pay gap/premium is reduced quite significantly once the sample is limited to employed workers, which is unsurprising. As shown in Table A1 and Table A2, there are significant differences in labor supply for LGB and straight workers. Gay men and straight women are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
There is a significant decrease in the hourly wage gap now and an insignificant decrease in the annual wage gap. There is also a significant decrease in the annual earnings premium for lesbian women but no change in their hourly wage premium. This result is consistent with the labor supply results in Table A1 and A2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
A.3 Occupation Fixed Effects
We show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table A4 and A5.
In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry. However, qualitatively the results seem to be similar to the main results in Table A1 and Table A3 . Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium. There is no longer a significant effect on the wage gap for gay men. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
Appendix Tables
Table A1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
Panel A: Men
(1) (2) (3) (4)
Labor Force Employed Labor Force Employed
Laws*SSP 0.013*** 0.014** 0.007 0.005
(0.005) (0.005) (0.005) (0.006)
SSP -0.074*** -0.079*** -0.068*** -0.073***
(0.003) (0.004) (0.003) (0.004)
State Laws Only X X
Observations 6,287,441 6,287,441 6,287,441 6,287,441
R-squared 0.135 0.114 0.135 0.114
Panel B: Women
(1) (2) (3) (4)
Labor Force Employed Labor Force Employed
Laws*SSP -0.013** -$0$016*** -0.013**
(0.006) (0.006) (0.006) (0.006)
SSP 0.081*** 0.073*** 0.080*** 0.072***
(0.004) (0.005) (0.004) (0.004)
State Laws Only X X
Observations 6,569,373 6,569,373 6,569,373 6,569,373
R-squared 0.091 0.088 0.091 0.088
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
Panel A: Men
(1) (2) (3) (4)
Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Laws*SSP 0.273* 0.089 0.180 -0.0744
(0.160) (0.080) (0.164) (0.0820)
SSP -2.196*** -0.575*** -2.106*** -0.477***
(0.114) (0.056) (0.0923) (0.0480)
State Laws Only X X
Observations 5,244,258 5,244,258 5,244,258 5,244,258
R-squared 0.034 0.011 0.034 0.011
Panel B: Women
(1) (2) (3) (4)
Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Laws*SSP -0.305** -0.184** -0.213 -0.171**
(0.136) (0.084) (0.134) (0.0790)
SSP 2.792*** 0.604*** 2.706*** 0.574***
(0.105) (0.066) (0.0924) (0.0577)
State Laws Only X X
Observations 4,366,603 4,366,603 4,366,603 4,366,603
R-squared 0.038 0.015 0.038 0.015
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A3: Effect of Anti-Discrimination Laws: Wages and Earnings
Panel A: Men
(1) (2) (3) (4) (5) (6) (7) (8)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP 0.028* 0.059 0.011 -0.041 0.0139 -0.00404 0.00639 -0.0581
(0.017) (0.050) (0.015) (0.035) (0.0184) (0.0525) (0.0168) (0.0373)
SSP -0.244*** -0.757*** -0.082*** -0.114*** -0.233*** -0.716*** -$0**$112***
(0.012) (0.038) (0.009) (0.023) (0.0108) (0.0354) (0.00745) (0.0184)
Employed Only X X X X
State Laws Only X X X X
Observations 6,287,441 6,287,441 5,244,258 5,244,258 5,244,258 6,287,441 5,244,258 6,287,441
R-squared 0.157 0.121 0.127 0.041 0.127 0.157 0.041 0.121
Panel B: Women
(1) (2) (3) (4) (5) (6) (7) (8)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP -0.019 -0.129** -0.024** -0.062** -0.0203 -0.153** -0.0144 -0.0453
(0.020) (0.066) (0.011) (0.029) (0.0202) (0.0668) (0.0112) (0.0297)
SSP 0.220*** 0.815*** 0.053*** 0.149*** 0.218*** 0.809*** 0.0448*** 0.132***
(0.014) (0.049) (0.007) (0.021) (0.0116) (0.0409) (0.00635) (0.0180)
Employed Only X X X X
State Laws Only X X X X
Observations 6,569,373 6,569,373 4,366,603 4,366,603 6,569,373 6,569,373 4,366,603 4,366,603
R-squared 0.136 0.106 0.123 0.036 0.136 0.106 0.123 0.036
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. Columns (3) and (4) are limited to those who are employed. All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Columns (5) - (8) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
Men Women
(1) (2) (1) (2)
Labor Force Employed Labor Force Employed
Laws*SSP 0.009*** 0.009** -0.009*** -0.009***
(0.003) (0.004) (0.003) (0.004)
SSP -0.037*** -0.045*** 0.039*** 0.036***
(0.002) (0.003) (0.002) (0.003)
Observations 6,287,441 6,287,441 6,569,373 6,569,373
R-squared 0.372 0.461 0.517 0.461
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients show the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients give the labor supply gap or premium. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A5: Anti-Discrimination Laws on Pay with Occupation FEs
Panel A: Men
(1) (2) (3) (4)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP 0.013 0.018 0.011 -0.028
(0.013) (0.035) (0.013) (0.033)
SSP -0.113*** -0.351*** -0.043*** -0.051**
(0.008) (0.025) (0.008) (0.022)
Employed only X X
Observations 6,287,441 6,287,441 5,244,258 5,244,258
R-squared 0.372 0.372 0.248 0.169
Panel B: Women
(1) (2) (3) (4)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP -0.015 -0.104*** -0.024*** -0.063**
(0.011) (0.034) (0.009) (0.025)
SSP 0.117*** 0.437*** 0.056*** 0.172***
(0.008) (0.028) (0.007) (0.018)
Employed only X X
Observations 6,569,373 6,569,373 4,366,603 4,366,603
R-squared 0.513 0.527 0.303 0.233
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A6: Sorting and Reporting
(1) (2) (3)
All SSP Male SSP Female SSP
Laws 0.000170 0.000633 -0.000304
(0.000489) (0.000572) (0.000573)
Observations 12,872,572 6,295,028 6,577,544
R-squared 0.068 0.072 0.065
  • Notes: Data comes from the 2005-2016 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law. The first column looks at both men and women with the next two columns separating the sexes. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A7: Differential Responses between Male and Female Same-Sex Partnerships
(1) (2) (3) (4)
One Earner HH Diff in Hours Worked Any Children Number of Children
Laws*FemSSP -0.00308 0.947*** 0.0344*** 0.0544***
(0.00868) (0.364) (0.00846) (0.0195)
FemSSP 0.00445 -0.729** 0.123*** 0.193***
(0.00717) (0.317) (0.00743) (0.0176)
Observations 73,181 73,181 73,181 73,181
R-squared 0.046 0.043 0.122 0.111
  • Notes: Data comes from the 2005-2016 yearly ACS comparing women in same-sex partnerships to men in same-sex partnerships. The first column examines if the household has only one-earner. The second column tests the difference in absolute terms of hours worked between the two partners. The third and fourth columns examines how households differ with having children. The first row of coefficients show the differential effect of anti-discrimination between lesbian and gay households. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.