### v2

Effect of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentialsthanks: We would like to thank Sandra Black, Rich Murphy, Steve Trejo, Gerald Oettinger, Dan Hamermesh, Kitt Carpenter, Dario Sansone, Chris Jepsen, Luis Faundez, and all the seminar participants of the University of Texas at Austin. No financial support was given to support this research. Correspondence: d.vamossy@pitt.edu.
Scott Delhommer University of Texas at Austin; sdelhommer@utexas.edu    Domonkos F. Vamossy University of Pittsburgh, d.vamossy@pitt.edu
(April 1, 2025)
This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on state and local anti-discrimination laws. Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws not only significantly narrow gaps in labor force participation and employment for male same-sex couples relative to men in different-sex couples, but also boost their percentile rank in the wage distribution. Our analysis reveals mostly null effects for female same-sex couples; however, in metropolitan areas these laws significantly reduce their employment compared to women in different-sex couples. One explanation for the reduced labor supply is that female same-sex couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
JEL Codes: J31, J71, J78, K31
1 Introduction
On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals. Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT. Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020. As a result, laws extending protected class status based on sexual orientation varied between states and localities. The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act. This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner. There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics. The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers. [^2] For a review of the impacts of legal access to same-sex marriage, see Badgett et al. (2024). For a broader review of the economics related to sexual orientation and gender identity, see Badgett et al. (2023).
We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts333(Badgett (1995); Klawitter (2011); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al. (2003); Carpenter (2005); Black et al. (2007); Antecol et al. (2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)), Tilcsik et al. (2015). The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
There have been previous studies examining how these state and local laws affect the pay gap, but these studies have been limited in their ability to identify causal effects (Klawitter and Flatt (1998); Gates (2009)). Klawitter and Flatt (1998) compared people in same-sex partnerships to those in different-sex partnerships in the 1990 Census and found that same-sex couples have no significant difference in wages in places with anti-discrimination laws. Gates (2009) uses the 2000 Census and conducts a similar analysis. Gates (2009) finds that gay/bisexual men in places with anti-discrimination laws have a 3% wage premium over gay/bisexual men in places without these laws, and lesbian/bisexual women have a 2% wage premium over lesbian/bisexual women in places without these laws [^4] Tilcsik (2011) found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws. These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework. [^5] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008). Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.666 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences. Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design. Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape. Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context. Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2005 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S. census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples. The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years. This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples. Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances. Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment.
This paper is the first to examine both local and state-level sexual orientation anti-discrimination laws in a quasi-experimental design. This paper is also the first to analyze the effects of any local anti-discrimination laws on sex, race, or sexual orientation in a modern difference-in-difference framework [^7] Using state variation in anti-discrimination laws has also been analyzed to understand racial- and sex-based discrimination (Neumark and Stock (2006); (Donohue III and Heckman, 1991);((Margo, 1995); Goldin and Margo (1992)). However, these studies on sex and race anti-discrimination laws also omit local laws, potentially skewing their estimates..
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws significantly reduce the gap in labor force participation and employment of men in male same-sex couples by 3.8 percentage points (78%) and 3.7 percentage points (77%), respectively. The laws improve men in male same-sex couples relative position in the wage distribution, with their percentile rank in the hourly wage and annual earnings distributions improving on average by approximately 2.7 and 3 percentiles, respectively. By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws. Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons. [^8] In Appendix B A, we replicate the findings of past research that finds men in male same-sex couples have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings. We also replicate findings that women in female same-sex couples have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
We explore theories for the differing effects of anti-discrimination laws on men in male same-sex couples and women in female same-sex couples in the Discussion section. Reduced discrimination improves job security for sexual minorities (Badgett, 2006), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (Chiappori, 1992; Lundberg and Pollak, 1993). These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that female same-sex households have significantly more children than male same-sex households after the passage of an anti-discrimination law.
Finally, our last contribution is the use of polling data on the support for same-sex marriage to examine the relationship between anti-discrimination laws and state sentiment toward LGB workers. One may expect that sentiment toward LGB workers would increase right before the passage of the laws, creating a selection issue. Alternatively, the anti-discrimination laws may normalize being a sexual minority, and improve sentiments toward LGB workers. We collected every poll on same-sex marriage for each state and year from the Pew Research Center. We show that the passage of state-wide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage. The states that passed anti-discrimination laws had parallel pre-trends in support of same-sex marriage before the passage of the laws. Afterward, those states had a significant and persistent increase in their support for same-sex marriage. The observed increase in support following the enactment of the law, rather than before, implies that policy changes can indeed influence public opinion. This notion aligns with the findings of Sansone (2019), suggesting that legislative actions can lead to shifts in societal attitudes. This is in contrast to the thermostatic model proposed by Wlezien (1995), which posits that public sentiment drives policy changes rather than the other way around. Increased favorability toward LGB people may act in conjunction with greater job security to improve labor market outcomes for LGB Americans.
2 Data
A common issue in analyzing the pay disparities between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment. We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition. Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves. [^9] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009).. The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that different-sex roommates would misclassify as unmarried partners.
The comparisons that we make are between people in same-sex partnerships (SSP), defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex partnerships (DSP). This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers. It compares those who are in a same-sex relationship to those who are in a different-sex relationship. Importantly, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual individuals in different-sex relationships as members of the control group. Because bisexual workers consistently earn less than both heterosexual and gay or lesbian counterparts ((Mize, 2016)), this classification introduces a bias toward underestimating pay disparities. Additionally, similar misclassification may occur if gay or lesbian individuals are categorized as heterosexual. Thus, the estimated pay gaps presented here likely represent conservative estimates.
We use the 2005-2019 yearly ACS from (Ruggles et al., 2019). We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked101010“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked.. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. Building on earlier city-level analyses such as (Klawitter, 2011), this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together within a state.
We report the state, city, and county laws and the year they were enacted in Table 1. In Figure 1, we show which counties had sexual orientation anti-discrimination laws in 2005 and 2019, respectively. In our analysis, laws are coded based on their initial application to private employment and housing, focusing on broad protections rather than sector-specific ones. For instance, while Delaware extended protections to state employees in 1999, our coding reflects 2009, the year broader protections were enacted. This approach could overlook the incremental impact of earlier, narrower protections on societal attitudes. The same principle applies to local laws, where initial protections for public employees in some cities might precede more comprehensive legislation, potentially influencing early shifts in public sentiment towards sexual minorities. The enforcement mechanisms of anti-discrimination laws, such as fines, mandatory training, and public rebukes, serve as significant deterrents, encouraging the adoption of more inclusive employment policies. The rise in successful legal challenges against companies concerning discrimination highlights the real legal and financial consequences they encounter, reinforcing the laws’ effectiveness in promoting workplace equality (for Human Resource Management (2023)).
The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^11] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. are noticeably absent from extending protection to LGB workers. One potential concern in this analysis is the endogenous adoption of laws. The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws. The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias. It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes. We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018). The inclusion of county-year fixed effects remove the need for county controls, addressing the lack of local support data for same-sex marriage. We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), designed for analyzing staggered policy adoptions like ours. This method accurately estimates the effects of anti-discrimination laws, considering their staggered implementation times.
In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships. Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston. The LGB population varies from the heterosexual population in many dimensions including geography.
We present descriptive statistics showing the differences in the same-sex couple population compared to the different-sex couple population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
3 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time and between same-sex and different-sex couples lends itself to a saturated triple difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county. Formally, we estimate this equation on labor supply:
$Y_{ipct}$ $=\alpha\;+\;\beta_{1}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr{]}\;+\;\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ (1)
$\quad+\theta_{ct}+\mu_{pc}+\kappa_{pt}+X_{ipct}^{\prime}\,\delta+\varepsilon_{ipct}.$
In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$, in partnership type $p$, residing in county $c$, at time $t$. These outcomes include labor force participation, employment status, and the percentile rank of real wages (in 2019 dollars). The binary variable $\mathrm{SSP}_{ipct}$ indicates whether an individual is in a same‐sex partnership (1) or a different‐sex partnership (0). $\mathrm{SSM}_{st}$ indicates whether state $s$ recognizes same‐sex marriage in year $t$, and $\mathbf{1}_{\{\mathrm{state}(c)=s\}}=1$ if county $c$ lies in state $s$. The error term $\epsilon_{ipct}$ is clustered at the county level to account for within-county correlation. [^12] The ACS only identifies metro counties, so the county fixed effect is close to a pseudo-county fixed effect. Rural counties within a state are compared together with a true county fixed effect for metro area counties. Our key variable of interest is the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$. [^13] Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter.
We include three sets of high‐dimensional fixed effects to account for a wide range of unobserved factors. First, $\theta_{ct}$ (county×$\times$×year) absorbs any time‐varying shocks or policy changes common to all individuals within a county in a given year. Second, $\mu_{pc}$ (partnership×$\times$×county) controls for any unobserved, time‐invariant characteristics that are specific to each partnership type within a county. Third, $\kappa_{pt}$ (partnership×$\times$×year) accounts for year‐specific shocks that affect each partnership type differentially. The vector $X_{ipct}^{\prime}\delta$ includes additional demographic controls (e.g., race, age, and education). The coefficient of interest, $\beta_{1}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same‐sex partners, net of any confounding from same‐sex marriage legalization. The summation term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\bigr{]}$ partials out state‐specific labor‐market impacts of marriage legalization on same‐sex couples.
The key variation used to estimate $\beta_{1}$ is the differential timing of when counties adopted ENDAs, and how this interacted with the labor market outcomes for individuals in same-sex partnerships relative to those in different-sex partnerships. This variation allows us to isolate the effect of anti-discrimination laws on labor market outcomes by comparing same-sex and different-sex couples across counties with and without ENDAs, before and after their passage. Thus, $\beta_{1}$ captures the differential impact of ENDA on labor supply and wage outcomes specifically for individuals in same-sex partnerships. [^14] We recognize the heterogeneous impact (e.g., Burn (2018) of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
We limit our sample to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. Following Chen and Roth (2024), we compute the percentile rank of annual and hourly real wages. [^15] Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. (2020), this extended timeframe might not precisely align with the implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.
To interpret $\beta_{1}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors are driving changes in the same-sex versus different-sex wage differential in states that do and do not have LGBT employment protections. One concern is that the distribution of workers in same-sex partnerships is not random, and states/counties that pass anti-discrimination laws have a higher concentration of workers in same-sex partnerships. However, with county-year fixed effects, comparisons are made among individuals within the same county in the same year, ensuring that workers in San Francisco are compared to others in San Francisco in that same year. A potential concern would be if there is sorting of workers in same-sex partnerships to areas with anti-discrimination laws protecting them, changing the composition of the area. To account for this, we remove individuals who moved in the past year as a robustness check. [^16] While we address potential sorting by excluding recent movers, unobserved migration patterns could still affect the composition of same-sex partnerships in jurisdictions with anti-discrimination laws.
Another potential concern is that same-sex marriage legalization may confound our estimates. We address this by explicitly partialing out any state-specific effects of marriage legalization for same-sex partners. In particular, we include the term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$, which absorbs any differential labor-market impacts associated with the varying timing of same-sex marriage legalizations across states. This ensures that our key coefficient, $\beta_{1}$, isolates the independent effect of the local Employment Non-Discrimination Act. While these controls mitigate potential bias, we caution that, despite our efforts, overlapping policies may still exert residual confounding influences on the estimated effects.
To address potential biases from staggered policy adoption, we employ the methodology of Borusyak et al. (2021). Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their approach uses only untreated observations to infer counterfactual outcomes for treated units. This strategy minimizes contamination and yields a more reliable estimate of $\beta_{1}$. Our identification hinges on the parallel trends assumption, that in the absence of anti-discrimination laws, labor market trends in counties with and without such laws would have evolved similarly. [^17] The difference-in-differences strategy hinges on the parallel trends assumption across groups and periods, with no anticipation of the laws prior to their implementation. We assess this assumption using event studies that estimate the dynamic effects of the policy, allowing us to observe pre-treatment trends and validate the plausibility of our identification strategy.
We report the coefficients estimated via Borusyak et al. (2021) in the main text using a fully saturated triple-difference (DDD) specification, and standard TWFE along with other difference-in-differences (DiD) estimators in Section 5. As we show in Section 5, the estimates for same-sex and different-sex individuals estimated separately are similar across estimators.
4 Primary Findings
4.1 Extensive & Intensive Margins
In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex partnerships. The coefficient $\beta_{1}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex partnerships and for women in female same-sex partnerships.
Panel A focuses on men in male same‐sex partnerships. Columns (1) and (2) show that anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77% respectively. These coefficients are statistically significant at the 1% level. Columns (3) through (6) address the intensive margin, indicating modest increases in weekly hours and weeks worked among all men. However, when the sample is restricted to employed men (Columns (5) and (6)), these effects become statistically insignificant. This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment.
Panel B reports analogous estimates for women in female same‐sex partnerships. In contrast, the extensive margin results for women (Columns (1) and (2)) yield statistically insignificant coefficients of $-0.37$ and $-0.50$ percentage points. Women in same‐sex partnerships exhibit higher labor force participation rates than women in different‐sex partnerships. This higher baseline may partially explain the lack of significant improvement on the extensive margin, as there is less room for gains. On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level) for all women, while Column (4) finds a small, statistically insignificant reduction in weeks worked. When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours (at the 5% level) alongside a significant decrease of 0.89 weeks worked (at the 1% level). These mixed findings suggest that anti-discrimination laws may lead employed women to work longer hours per week but in fewer weeks overall, reflecting a complex adjustment on the intensive margin.
4.2 Wages
As shown in Table 4, anti-discrimination laws significantly affect labor force participation and employment outcomes, thereby potentially altering the composition of the employed sample and raising endogeneity concerns for wage analyses. Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women. As Chen and Roth (2024) argue, this robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex partnerships.
For men in same-sex partnerships, the increased employment rates reported in Table 4 are accompanied by significant wage gains. Specifically, anti-discrimination laws boost hourly wages by roughly 2.7 percentiles and annual earnings by 3 percentiles on average. Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.6 and 3.4 percentage points, respectively, although the effect at the 75th percentile is not significant. In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 5.2, and 3.7 percentage points at the 25th, 50th, and 75th percentiles, respectively. For women in same-sex partnerships, the wage distribution remains largely unaffected by anti-discrimination laws, as evidenced by statistically insignificant estimates across the board. This finding is consistent with the muted effects on employment outcomes observed in Table 4 for women.
These findings suggest that the wage gains for men in same-sex partnerships operate primarily through increased employment. New entrants to the labor force, individuals who were previously at the bottom of the wage distribution, enter at various points, thereby shifting the overall distribution upward.
4.3 Impact Dynamics
We estimate Equation (1) using event study regressions following Borusyak et al. (2021) to trace the evolution of the impact estimates. Figure 3 displays the results for extensive labor supply outcomes: Panel (a) shows labor force participation rates, while Panel (b) presents employment levels. For men in male same‐sex partnerships, both panels reveal an immediate increase in labor force participation and employment following the enactment of ENDAs, although this positive effect wanes over time. In contrast, we see no impacts on women in same-sex partnerships. The pre-treatment trends are largely parallel, supporting the plausibility of the parallel trends assumption, yet the post-treatment dynamics diverge sharply by gender. The evidence here suggests that ENDAs lead to a temporary boost in labor market entry for men in same-sex partnerships, while not affecting employment and participation rates for women in same-sex partnerships.
Figure 4 presents the event study regressions for male and female pay disparities. The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex partnerships. However, the difference between the one-year and four-year post‑enactment estimates is relatively modest, suggesting that the initial wage improvement for men stabilizes and remains persistent over time. We find no statistically significant effects on wage disparities for women in any post‑ENDA period. These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex partnerships, the impact on wage outcomes for women in same-sex partnerships is insignificant.
5 Robustness Checks
To ensure the robustness of our findings on the labor market impacts of anti-discrimination laws, we conduct a series of sensitivity analyses addressing potential sources of bias. Specifically, we examine (i) the endogeneity of law adoption using state-level support for same-sex marriage as a proxy for public sentiment, (ii) the potential confounding influence of same-sex marriage legalization, (iii) the role of migration patterns, (iv) geographic heterogeneity between metropolitan and non-metropolitan areas, (v) reporting changes associated with ACS redesign, (vi) the impact of applying sample weights, and (vii) the validity of our identification strategy using placebo tests and alternative estimators.
5.1 Endogenous Adoption of Anti-Discrimination Laws
Anti-discrimination laws are not randomly distributed. Locally, anti-discrimination laws are concentrated in larger cities, and state laws are to be concentrated in more liberal states that presumably are more accepting of and more favorable to LGB workers. The main concern is that there is an unobservable factor like general sentiment toward LGB workers that affects both the passage of laws as well as the labor market outcomes for LGB workers. In our main regressions, we alleviate this concern by using county-year fixed effects.
Nonetheless, we investigate the endogenous adoption of anti-discrimination laws at the state level. Specifically, we use state-level polling information on support for same-sex marriage as a proxy for general sentiment toward LGB workers. It is possible that controlling for state-level polling is not the best way to capture sentiment toward LGB workers since it is possible to discriminate against people based on sexual orientation and still support their right to marry. However, it seems plausible that the changes in state-level support for same-sex marriage are highly correlated with changes in sentiment toward LGB workers such that it will suffice for a suitable proxy. To better get at the question of endogenous adoption of laws, we create an event-study plot showing how state laws change support for same-sex marriages. Specifically, we estimate this equation:
$\text{Support}_{st}$ $=\sum_{\begin{subarray}{c}j=-4\\ j\neq-4\end{subarray}}^{4}\psi_{j}\,\mathbbm{1}\Bigl{(}\text{ENDA}_{st}=j\Bigr{)}+\phi_{s}+\delta_{t}+\epsilon_{ts}.$ (2)
Supportst denotes the level of support for same-sex marriage in state $s$ during year $t$, while $\mathbbm{1}(\text{ENDA}_{st}=j)$ is an indicator variable equal to 1 if state $s$ has had a sexual orientation anti-discrimination law for $j$ years by time $t$. We estimate the coefficients $\psi_{j}$, which capture the effect of anti-discrimination laws on support for same-sex marriage, using an event study framework within a difference-in-differences design. This strategy relies on the identifying assumption that, in the absence of anti-discrimination laws, support for same-sex marriage in states that passed these laws would have followed parallel trends to those in states that did not. Figure 5 presents the event study plot.
The event study indicates that state-level anti-discrimination laws significantly boost support for same-sex marriage. Prior to the legislation, trends in support moved in parallel across states regardless of their legal stance, suggesting that, absent the laws, these trajectories would have remained similar. Although one might argue that the passage of such laws was driven by a pre-existing increase in support for same-sex marriage, an endogenous explanation would predict a gradual rise in support leading up to the law’s enactment, with little change at the moment of passage. Instead, we observe a distinct, significant jump in support coinciding with the implementation of the legislation. This pattern implies that the legislative changes themselves are likely driving the increase in support for same-sex marriage, and possibly improving overall attitudes toward LGB individuals.
Moreover, the observed jump in state-level sentiment toward LGB workers following the passage of anti-discrimination laws may reflect an unobserved corresponding increase in local sentiment. This local-level change could act as a mechanism through which these laws impact the labor supply decisions of LGB workers.
5.2 Marriage Laws
Same-sex marriage legalization may confound our identification strategy. For example, Utah legalized same-sex marriage in 2015, the same year it passed its statewide sexual orientation anti-discrimination bill, while Iowa and Arizona experienced similar overlaps (with Iowa enacting its anti-discrimination law in 2007 and legalizing same-sex marriage in 2009, and Arizona in 2014). Prior studies have shown that same-sex marriage legalization can differentially affect employment outcomes for same-sex couples. To address this issue, we included fixed effects that interact same-sex couple status with state-level same-sex marriage legalization. Nevertheless, to further isolate the impact of anti-discrimination laws, we restrict our sample to periods before same-sex marriage became legal in each state (e.g., excluding observations from 2008 onward for states that legalized in 2008).
Table 6 presents results restricting the sample to periods before same-sex marriage legalization. Although this approach removes nearly half of the same-sex partnership observations and reduces statistical power, the qualitative patterns remain robust. Specifically, the point estimates for both labor force participation and employment remain comparable, exceeding 3 percentage points, with labor force participation retaining statistical significance. However, the employment coefficient for men declines from 0.037 to 0.0319, becoming borderline insignificant (p-value = 0.102), and similar shrinkages are observed on the intensive margin. In contrast, for women in same-sex partnerships, the intensive margin estimates are slightly larger. Overall, this sensitivity analysis indicates that our identification strategy is not driven by concurrent same-sex marriage policies, as the observed effects of anti-discrimination laws on labor market outcomes for same-sex couples persist despite the reduced sample size.
5.3 Migration
Next, we account for the possibility that migration patterns could influence our results. If LGB individuals were disproportionately moving to areas with recently enacted anti‑discrimination laws, it could bias the treatment effects. To address this concern, we restrict our sample to individuals who had not moved in the past year and re‑estimate the models presented in Table 4. By focusing on a more stable population, we can better isolate the effects of anti‑discrimination laws on labor outcomes. The corresponding estimates, reported in Table 7, remain consistent with the full sample results, providing further robustness and suggesting that migration does not significantly alter our conclusions. Interestingly, the estimates for men in same‑sex partnerships are slightly larger for this subset, which may indicate that the positive employment effects are even more pronounced among those who are less mobile.
5.4 Metropolitan Areas
We estimate our model specifically for residents in metropolitan areas to address concerns about potential differences between urban and rural populations. By focusing on metro areas, where labor markets may function differently due to higher population density and greater economic diversity, we aim to test whether the effects of anti-discrimination laws vary by geographic location.
Table 8 shows that when the sample is restricted to metropolitan areas, the overall patterns for men in same‑sex partnerships remain consistent, though the estimated effects of anti‑discrimination laws on labor supply are slightly smaller than in the full sample. This may suggest that the urban labor market environment, with its higher baseline employment levels and different institutional dynamics, somewhat tempers the positive impact seen elsewhere. In contrast, for women in same‑sex partnerships, the estimates become significantly negative across several outcomes, labor force participation, employment, and weeks worked. One plausible explanation is that ENDAs may improve employment stability in urban labor markets, and with high childcare costs in metro areas, a stable job by one partner might encourage women to opt out of the workforce. With a secure income from the other partner, some women may choose to stay home to manage childcare responsibilities rather than work additional hours or remain continuously employed. This pattern highlights clear geographic heterogeneity, as the dynamics in urban settings differ markedly from those in the broader sample for women in same‑sex partnerships.
5.5 Reporting Changes
Prior research has highlighted higher rates of misclassification errors before 2008, where different-sex couples were sometimes incorrectly identified as same-sex couples due to typos in the sex variable. In response, and considering the 2008 ACS redesign aimed at reducing such errors (O’Connell and Lofquist (2009)), we restrict our sample to 2008–2019 to test the robustness of our findings reported in Table 4. Table 9 presents the estimates based on this restricted sample.
For men in same-sex partnerships, the results remain qualitatively similar to our full sample analysis. The estimated impacts on both labor force participation and employment continue to be positive and statistically significant, with coefficients that are comparable in magnitude to those in Table 4. Among women, the extensive margin estimates remain small and statistically insignificant, while the intensive margin results are largely consistent with the full sample, albeit with some variations in magnitude.
These findings suggest that potential misclassification errors in earlier years do not drive our main conclusions. The consistency of the results between the full sample (2005–2019) and the restricted sample (2008–2019) reinforces the robustness of our empirical strategy and the validity of our conclusions regarding the impact of anti-discrimination laws on labor market outcomes for same-sex couples.
5.6 Weighting
Although the ACS recommends applying sample weights to ensure representativeness, we conduct a robustness check by re‐estimating our models without these weights, following Solon et al. (2015). This approach allows us to verify that our main results are not driven solely by the weighting scheme.
The results in Table 10 indicate that the overall impacts of anti‑discrimination laws on labor supply remain robust even when ACS weights are removed. For men in same‑sex partnerships, the magnitude of the estimated effects on both labor force participation and employment is slightly smaller compared to the weighted estimates, but they continue to be statistically significant. This suggests that the positive impact of these laws for men is not driven solely by the weighting procedure. For women, the pattern of the results is similar to the weighted sample, with some variations in magnitude across different labor outcomes.
5.7 Placebo Tests and Alternative Estimators
We next conduct a series of placebo tests and apply alternative estimators to assess the impact of anti‑discrimination laws on labor market outcomes among both same‑sex and different‑sex couples. Specifically, we utilize four distinct methods to assess the impact of anti‑discrimination laws on labor market outcomes among same‑sex and different‑sex couples: TWFE, Callaway and Sant’Anna (2021), Sun and Abraham (2021), and Borusyak et al. (2021). As a placebo test, we apply the DiD framework to different‑sex couples, who are not the target of these laws, and thus we expect no significant changes in their labor market outcomes. Panel (a) of Tables 11 and 12 report the DiD estimates for different‑sex couples, while Panels (a) and (c) of Figures 6 and 7 display the corresponding event study plots.
Our placebo tests show that the DiD estimates yield no statistically significant impacts on labor force participation or employment rates for different‑sex couples. The pre‑ENDA coefficients are insignificant, confirming the absence of differential trends prior to the law’s enactment. Most post‑ENDA coefficients remain statistically insignificant, with only a few minor exceptions of small magnitude, largely consistent with a 5% false positive rate.
Next, Panel (b) of Tables 11 and 12 presents the impact of anti‑discrimination laws on same‑sex couples, while Panels (b) and (d) of Figures 6 and 7 display the corresponding event studies. In these analyses, we compare labor market outcomes for same‑sex couples in regions that have enacted the laws with those in regions that have not. The results reveal that male same‑sex couples in areas with anti‑discrimination laws experience significant improvements in labor market outcomes, particularly through increased labor force participation and higher employment rates. For female same‑sex couples, however, we observe no significant effects on the extensive margin; instead, the positive and significant impact is confined to the intensive margin, as evidenced by an increase in average weekly hours worked. Hence, our alternative estimators yield similar point estimates and comparable confidence intervals, reinforcing the robustness of our findings.
6 Discussion
6.1 Differences in Response by Sex
The results reveal that anti-discrimination laws affect household labor allocation differently for same-sex couples by gender. Men in male same-sex partnerships experience a reduction in their labor supply gap relative to men in different-sex partnerships, while there are no significant average effects for women in female same-sex partnerships.
One potential explanation is that changes in job security, stemming from reduced discrimination, alter the household’s allocation of labor between market and non‑market activities. As documented by Badgett (2006), reduced discrimination in the labor market improves employment prospects and job security for sexual minorities. With greater job security for the primary earner, the secondary earner may reduce market participation in favor of increased household production, including child rearing. This pattern of adjustment is consistent with models of household bargaining and collective decision‑making (see, e.g., Chiappori (1992) and Lundberg and Pollak (1993)), which predict that improvements in labor market conditions for one partner can induce a reallocation of tasks within the household. In this framework, the observed increase in intra‑household differences in hours worked may reflect a shift toward greater specialization, with one partner focusing more on market production while the other devotes more time to household responsibilities. Thus, the differential responses between male and female same‑sex partnerships could arise from distinct household preferences and constraints.
Households consisting of women in same-sex partnerships typically have more children than those consisting of men in same-sex partnerships. Given this difference, it would be unsurprising to see female same-sex households shift more toward a specialization of household labor compared to male same-sex households. This difference in specialization could help explain why anti-discrimination laws have different effects on women in same-sex partnerships and men in same-sex partnerships.
To test the different responses by men and women in same-sex partnerships to anti-discrimination laws, we implement an alternative difference-in-difference model. Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level. Do women in same-sex partnerships change their household division of labor relative to men in same-sex partnerships following anti-discrimination laws? To answer this question, we collapse the data to the household level and examine households instead of individuals. Specifically, we estimate this model:
$Y_{jct}$ $=\alpha_{0}\,ENDA_{ct}+\alpha_{1}\,FemSSP_{j}+\alpha_{2}\,\bigl{(}ENDA_{ct}\times FemSSP_{j}\bigr{)}+\mu_{c}+\delta_{t}+\epsilon_{jct}$ (3)
The notation is the same as before with two main differences. First, our interaction term uses an indicator for whether household $j$ is a female same‐sex partnership, denoted by $FemSSP_{j}$. Second, $Y_{jct}$ now represents outcomes measured at the household level. The coefficient $\alpha_{2}$ is our parameter of interest and captures how female same‐sex households differ in their response to anti‑discrimination laws relative to male same‑sex households. The first outcome examined is an indicator for whether the household is a one‑earner family. We also analyze the intra‑household division of labor by measuring the absolute difference in weekly hours worked between partners, thereby capturing both the extensive and intensive margins of labor supply. Finally, we explore the potential effects on child rearing by examining both the presence of children in the household and the number of children present.
We present the results in Table 13. Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin. There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for male same-sex and female same-sex households. An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership. Female same-sex households do see a significant increase in the likelihood of having any child and the number of children relative to male same-sex households.
This evidence suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws. Female same-sex households begin to have more children than male same-sex households. It may appear in the main results that anti-discrimination laws do not help women in female same-sex partnerships. However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security. The corresponding changes in the labor market could represent an increase in welfare for female same-sex households if they value their new children more than their lost hours worked and subsequently lost wages. Consistent with the theory and empirical findings, female same-sex households could become more secure with one woman being the primary earner following anti-discrimination laws.
6.2 Mechanisms
The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for men in male same-sex partnerships. It is possible that these laws are binding and eliminate significant discrimination in the labor market. However, it is relatively easy for an employer to create a reason to fire an employee rather than firing someone specifically for being gay. It is also difficult to accurately determine someone’s sexual orientation by looking at them, unlike race and sex. Although it may be easier to determine someone’s sexual orientation conditional on them being in a same-sex relationship. Sexual orientation anti-discrimination laws may not be binding in a meaningful way that eliminates discrimination.
One mechanism that the laws could be affecting labor market outcomes is through increased positive sentiment toward LGB workers. We show in Figure 5 that following state anti-discrimination laws there is a significant and persistent increase in the percentage of people in that state that support same-sex marriage. Our regression results already condition on the change in the county-level support for same-sex marriage via fixed effects. It seems plausible that local law changes could significantly change local sentiment and impact the labor supply and pay of LGB workers. Changing public sentiment in conjunction with greater protection for LGB workers seems more plausible as the mechanism than simply greater protection in the workplace.
6.3 Threats to External Validity
One limitation of this research is that we are unable to view all LGB workers. The analysis is specifically conducted on individuals in same-sex partnerships and different-sex partnerships. The data does not allow for the identification of single LGB workers, whose outcomes may be significantly different than LGB workers in partnerships. Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership. Bisexual individuals represent a larger group than lesbian or gay individuals ((Gates, 2011)), and our specification erroneously classifies bisexual people in different-sex relationships as heterosexual. This misclassification biases estimates toward zero, as bisexual workers face significant wage penalties ((Mize, 2016)) and their inclusion in the reference group depresses the heterosexual average. Future work must disentangle these groups to avoid obscuring inequities.
Another potential concern in using partnerships to infer sexual orientation is that people in a same-sex partnership may be the group that is most at risk of discrimination. It seems likely that single LGB workers can more plausibly stay in the closet with their co-workers compared to their counterparts in a same-sex partnership. This analysis may be capturing the effect of anti-discrimination laws on the group that is most likely to be affected. These results may have a larger effect size than what one would find in examining the generalized LGB population.
These biases are for the broader LGB population and external validity. Notably, these potential biases do not affect internal validity. Translating the effect of anti-discrimination laws for people in same-sex partnerships to all other LGB workers is not immediately obvious. It is dependent on several factors that are unknowable in this analysis such as the wages and labor supply for single LGB workers and bisexual workers in different-sex partnerships and how discrimination affects LGB people for those in same-sex partnerships.
Furthermore, it is important to emphasize that this study cannot identify transgender individuals in the ACS, nor can it estimate the impact of anti-discrimination laws specifically protecting gender minorities. This restricts the ability to generalize findings to the entire LGBTQ+ community.
7 Conclusion
This analysis is the first quasi-experimental study examining the impact of both local and state anti-discrimination laws on labor supply and wage disparities between workers in same-sex and different-sex partnerships. Utilizing a novel city-level panel dataset capturing the granular detail of anti-discrimination laws, we find that these laws significantly reduce sexual orientation-based inequalities in labor market outcomes. Specifically, anti-discrimination laws decrease the labor force participation difference by 3.8 percentage points (78%) and employment difference by 3.7 percentage points (77%) for men in male same-sex partnerships, with insignificant effect for women in female same-sex partnerships. We also document important wage effects. Sexual orientation anti-discrimination laws lead to significant wage gains for men in same-sex partnerships, with their percentile rank in the hourly wage and annual earnings distribution improving on average by approximately 2.7 and 3 percentiles, respectively. Moreover, these laws increase the probability of ranking above the 25th percentile in hourly wages by 3.6 percentage points and above the 50th percentile by 3.4 percentage points, with similar improvements observed in the annual earnings distribution. By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws.
We identify increased positive societal sentiments toward LGB Americans, reflected in growing support for same-sex marriage, as a potential mechanism through which these laws improve labor market outcomes. Furthermore, we explain differential responses to these laws between male and female same-sex households using Becker’s household specialization model and provide empirical support for this explanation.
Prior to the landmark Supreme Court decision in Bostock v. Clayton County (2020), many states permitted employers to fire employees based solely on sexual orientation. While Bostock extended federal protections under Title VII to LGBTQI+ workers, legislative momentum for anti-discrimination laws at the state level has slowed significantly since then. The Equality Act, reintroduced in the 118th Congress on June 21, 2023, seeks to expand federal civil rights laws to explicitly include sexual orientation and gender identity. As of March 2025, the bill is pending in the Senate, facing significant opposition from some senators, particularly concerning its implications for religious freedoms and women’s rights in certain contexts. At the state level, legislative momentum for anti-discrimination laws has slowed. Between 2009 and 2019, more states barred cities from protecting their residents from discrimination than those that expanded protections. Additionally, recent developments have seen states like Arkansas and Tennessee actively prohibiting localities from enacting their own anti-discrimination measures.
Our research provides critical and timely insights into the impact of anti-discrimination laws on labor market outcomes for individuals in same-sex partnerships. By offering a comprehensive evaluation of these laws’ effects, our findings meaningfully inform ongoing policy discussions and judicial decisions, including deliberations by the Supreme Court of the United States, about extending employment protections to the LGB community.
Appendix A Additional Robustness Checks
A.1 Event Studies
Figures A1, A2, A3, A4, A5 present the event study regressions corresponding to the robustness checks discussed in Sections 5.2, 5.3, 5.4, 5.5, 5.6. In each figure, we trace the evolution of labor supply impacts on the external margin relative to the enactment of anti-discrimination laws. Across all specifications, the pre-treatment coefficients are statistically insignificant, and the post-treatment dynamics closely mirror those observed in our main analyses. Specifically, estimates for women in metropolitan areas are smaller on average, unweighted models and those estimated before the legalization of same-sex marriage typically produce slightly lower coefficients, and both the post-2008 sample and non-movers yield somewhat larger effects.
A.2 Same-Sex Marriage Fixed Effects
The results in Table A1 reinforce the robustness of our primary findings for men. Although the extensive margin effects for men in same-sex partnerships are slightly lower (0.0267 for labor force participation and 0.0270 for employment) than in our primary specification, they remain statistically significant, with generally smaller intensive margin estimates. For women, the point estimates on the extensive margin are quite similar in Table A1 and Table 6. However, the employment effect for women turns negative and is marginally significant (–0.0212, p $<$ 0.10), which contrasts with the mostly null effects observed elsewhere, aside from the negative effects noted in metropolitan areas. This conflicting evidence suggests that further research is needed to better understand the impacts on women. Our hypothesis remains that high child-care costs in metropolitan areas may be an important contributing factor.
A.3 Occupation Fixed Effects
LGB workers tend to sort to different occupations and industries based on sex and sexual orientation [Black et al., 2007]. They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts. It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
Although controlling for occupation can pose selection concerns, because individuals not in the labor force typically lack occupation codes, we nonetheless examine within-occupation effects of anti-discrimination laws on wages and labor supply. Specifically, we employ regressions with four-digit occupation fixed effects to assess how wage and labor supply disparities shift following the enactment of these policies. [^18] In order to estimate within-occupation effects while still capturing individuals who are not in the labor force, we include a separate “no occupation” category for those without valid occupational codes. This approach allows us to retain non-employed respondents in the analysis, rather than excluding them outright due to missing occupation information. We acknowledge that categorizing all non-employed individuals into a single group can mask heterogeneity in reasons for not working (e.g., unemployment, home production, or retirement). Nonetheless, by creating a distinct “no occupation” category, we preserve a measure of occupational affiliation for those who are employed while ensuring that non-employed individuals remain part of the sample. One potential mechanism is that workers in male same-sex partnerships may transition into more in-demand occupations, thereby mitigating relative labor supply and wage gaps. However, as shown in Tables A2 and A3, even after accounting for occupation, men in same-sex partnerships exhibit significant changes in both labor supply and wages. These findings suggest that the impact of anti-discrimination laws is not merely a by-product of occupational sorting. Nonetheless, caution is warranted when interpreting these results, given the inherent limitations of conditioning on occupation in a broader labor force context.
Appendix Tables
Table A1: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Same-Sex Marriage Fixed Effects
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Panel A: Men
ENDA $\times$ SSP 0.0267∗∗ 0.0270∗∗ 0.5377 0.8852 -0.4480 -0.3068
(0.0129) (0.0133) (0.4982) (0.6173) (0.3629) (0.2818)
Observations 3,798,599 3,798,599 3,798,599 3,798,599 3,073,844 3,073,844
Mean of DV 0.8254 0.7903 35.5714 39.9865 42.4118 47.6751
Employed Only X X
Panel B: Women
ENDA $\times$ SSP -0.0130 -0.0212 0.8477 -0.7081 0.9893∗∗∗ -0.7657∗∗
(0.0108) (0.0115) (0.5163) (0.5760) (0.3081) (0.3166)
Observations 3,982,894 3,982,894 3,982,894 3,982,894 2,691,849 2,691,849
Mean of DV 0.8200 0.7846 34.5069 39.5937 41.2556 47.2540
Employed Only X X
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. The regressions are estimated separately by sex: Panel A for men, while Panel B for women. The coefficients indicate the impact of anti-discrimination laws. The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws. Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours. We exclude $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ from Equation (1) for this table. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A2: Anti-Discrimination Laws on Labor Supply with Occupation FEs
(1) (2) (3) (4) (5) (6)
Extensive Margin Intensive Margin
Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Panel A: Men
ENDA $\times$ SSP 0.0228∗∗ 0.0246∗∗ 0.5170 0.8743∗∗ -0.1462 -0.0550
(0.0093) (0.0100) (0.3720) (0.4334) (0.3307) (0.2797)
Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,073,797 3,073,797
Mean of DV 0.8254 0.7903 35.5714 39.9865 42.4118 47.6751
Employed Only X X
Panel B: Women
ENDA $\times$ SSP -0.0064 -0.0132 0.5614 -0.8344 0.5854∗∗ -0.9524∗∗∗
(0.0098) (0.0111) (0.4008) (0.4630) (0.2939) (0.3219)
Observations 3,982,816 3,982,816 3,982,816 3,982,816 2,691,783 2,691,783
Mean of DV 0.8200 0.7846 34.5069 39.5937 41.2556 47.2540
Employed Only X X
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. The regressions are estimated separately by sex: Panel A for men, while Panel B for women, and include four-digit occupation fixed effects. The coefficients indicate the impact of anti-discrimination laws. The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws. Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A3: Anti-Discrimination Laws on Pay with Occupation FEs
Hourly Wage Annual Earnings
(1) (2) (3) (4) (5) (6) (7) (8)
Percentile $>$25th $>$50th $>$75th Percentile $>$25th $>$50th $>$75th
Panel A: Men
ENDA $\times$ SSP 1.9105∗∗ 0.0241∗∗ 0.0262 0.0239 2.2973∗∗∗ 0.0241∗∗ 0.0445∗∗∗ 0.0340∗∗
(0.9218) (0.0113) (0.0147) (0.0147) (0.8353) (0.0113) (0.0148) (0.0132)
Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542
Mean of DV 52.0273 0.7917 0.5653 0.3121 52.4128 0.7917 0.5746 0.3163
Panel B: Women
ENDA $\times$ SSP -0.9428 -0.0023 -0.0232∗∗ -0.0096 -1.1544 -0.0023 -0.0217 -0.0106
(0.6914) (0.0097) (0.0116) (0.0148) (0.6491) (0.0097) (0.0118) (0.0154)
Observations 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816
Mean of DV 49.3832 0.8002 0.5246 0.2564 49.5588 0.8002 0.5324 0.2551
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment. The regressions are estimated separately by sex and include four-digit occupation fixed effects. The coefficients in the first row indicate the impact of anti-discrimination laws. Columns (1–4) report estimates for hourly real wage ranks, while Columns (5–8) report estimates for annual real wage ranks. In Columns 2–4 and 6–8, the estimates are indicator variables for being above the 25th, 50th, and 75th percentiles, respectively. All income variables are measured in 2019 dollars and expressed as percentile ranks. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Appendix Figures
((a)) Labor Force Participation
((b)) Employment
Figure A1: Labor Supply Impacts on the Extensive Margin: Same-Sex Marriage Legalization
Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws. Sample is limited to observations before same-sex marriage legalization. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level.
((a)) Labor Force Participation
((b)) Employment
Figure A2: Labor Supply Impacts on the Extensive Margin: Non-Movers
Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws. Sample is limited to individuals who have not moved in the past year. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level.
((a)) Labor Force Participation
((b)) Employment
Figure A3: Labor Supply Impacts on the Extensive Margin: Metropolitan Areas
Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws. Sample is limited to individuals who live in metropolitan areas. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level.
((a)) Labor Force Participation
((b)) Employment
Figure A4: Labor Supply Impacts on the Extensive Margin: 2008-2019
Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws. Sample is limited to cover 2008 through 2019. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level.
((a)) Labor Force Participation
((b)) Employment
Figure A5: Labor Supply Impacts on the Extensive Margin: Weighting Schemes
Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws. We compare estimates weighted by ACS weights and without. Coefficients with 95% confidence intervals. Standard errors are clustered at the county level.
Online Appendix*: Effects of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentials
In the original version of our paper, we employed a traditional difference-in-differences (DiD) setup, incorporating a range of controls to assess the impact of anti-discrimination legislation on labor market outcomes for LGB workers. This conventional approach served as the foundational analysis for our initial findings. To ensure transparency and thoroughness in our research, we are including these initial estimates in this section. Furthermore, in response to recent advancements in DiD methodologies, we use Borusyak et al. [2021] in the main text to estimate the impact of anti-discrimination laws. Our analysis reveals that the application of these modern DiD techniques does not substantially alter the point estimates we initially obtained. This consistency across methodological frameworks underscores the robustness of our findings, reinforcing the reliability of our conclusions regarding the effects of anti-discrimination laws on LGB labor market outcomes.
In the TWFE setup, we control for sentiment toward LGB workers in a given state at a given year by including polling information on support for same-sex marriages. Unfortunately, we cannot observe polling at a local level, as such the polling measure is an incomplete proxy for sentiment toward LGB workers. We collected every poll that Pew Research Center published from 2005-2016 to pull each poll regarding support for same-sex marriage. The 28 polls asking about same-sex marriage were aggregated to get the percentage of people that supported same-sex marriage by state by year. The polling information contains data on every state for every year, except for Alaska and Hawaii, which are missing polling information for 2005-2008.
O.4 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, workers in same-sex versus different-sex partnerships. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county.
Formally, we estimate this equation on labor supply:
$LS_{ict}=\alpha_{0}ENDA_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}ENDA_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
Where $LS_{ict}$ is the labor force participation or employment status for person $i$ in county $c$ in year $t$. $ENDA_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership. $X_{i}$ is a vector of person-specific covariates including race, children, along with age and education level - each represented up to a quadratic term to capture potential nonlinear effects. $\gamma_{st}$ is a vector for state $s$ covariates on the legality of same-sex marriage in year $t$ as well as polling for the support of same-sex marriage. It includes the percentage of people that “strongly support,” “support,” “oppose,” and “strongly oppose” same-sex marriage in a given state in a given year. Finally, $\mu_{c}$ gives county fixed effects, $\delta_{t}$ gives year fixed effects, and $\epsilon_{itcs}$ is the error term [^19] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties.. Our sample is limited to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. We cluster the standard errors at the county level, as that is the level of the treatment in this specification. [^20] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. [2020], there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The coefficients of interest are $\alpha_{2}$, which will give the effect of the law on the labor supply gap/premium, and $\alpha_{1}$, which gives the labor supply gap/premium in the absence of anti-discrimination law. Our specification assumes that the effect of the laws is the same across states and localities, and we will be estimating an average effect. We estimate a similar equation for wages:
$\ln(Y_{ict}+1)=\beta_{0}\text{ENDA}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{ENDA}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
Where $Y_{ict}$ is the real value, in 2005 dollars, of person $i$’s wages in county $c$ year $t$. The rest of the notation is the same as above. The coefficients of interest are $\beta_{2}$, which will give the effect of the law on the pay gap/premium, and $\beta_{1}$, which gives the pay gap/premium. [^21] The $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum, capturing nuances missed by focusing solely on employed workers.
O.5 TWFE Results
In Table O1, we present the regression on the extensive margin of labor supply, with Panel A showing the effect of anti-discrimination laws on the labor supply of men in male same-sex partnerships, and Panel B showing the effect for women in female same-sex partnerships. We present the full specification in columns (1) and (2) and include a specification without accounting for local anti-discrimination laws in columns (3) and (4) to show how failing to account for local laws would lead to incorrect inference.
Consistent with the previous literature, we find that women in female same-sex partnerships have a labor supply premium and men in male same-sex partnerships have a labor supply penalty on the extensive margin. Women in female same-sex partnerships are 7.3% more likely to be employed and men in male same-sex partnerships are 7.9% less likely to be employed than their different-sex counterparts. Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for men in male same-sex partnerships. The labor force participation rate gap is reduced by 1.3 p.p. (18%), and the employment gap is reduced by 1.4 p.p. (18%). Conversely, the labor force participation and employment premium that women in female same-sex partnerships have over their different-sex counterparts is reduced by 1.3 p.p. (16%) and 1 p.p. (14%), respectively. As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws using only state laws may lead to the erroneous conclusion that these laws have no significant effect on the labor force participation and employment gaps for men in male same-sex partnerships. However, the effect on the extensive margin of labor supply for women in female same-sex partnerships looks similar when estimating the effect of just state anti-discrimination laws.
In Table O2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. Similar to Table 4, we include the results when only estimating the effects of the state laws in columns (3) and (4) to show the contribution of including local laws in estimating the effect of anti-discrimination laws.
We replicate past findings where men in male same-sex partnerships supply significantly less labor and women in female same-sex partnerships supply significantly more labor on the intensive margin. As with Table O1, we find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for men in male same-sex partnerships as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for women in female same-sex partnerships is significantly reduced as well. The effect of anti-discrimination laws on weekly hours worked becomes insignificant for men in male same-sex partnerships and women in female same-sex partnerships when only estimating off of state-wide anti-discrimination laws.
In Table O3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a log-transformation, to handle zero values so the changes can be interpreted in percentage terms. We run the regressions separately by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table O1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers. We include results by only using state anti-discrimination laws, excluding local laws. We present those results in columns (5) - (8).
Men in male same-sex partnerships see a significant decline in the hourly wage gap by 2.8 p.p. (11%) at the 90% significance level and see an insignificant decline in the annual earnings gap by 5.8 p.p. On the other hand, women in female same-sex partnerships see a significant change in their annual earnings premium by 13 p.p. (16%) and an insignificant reduction in their hourly wage premium. Columns (3) and (4) replicate the past literature where employed men in male same-sex partnerships make about 8% to 11% less than their different-sex counterparts and employed women in female same-sex partnerships make about 5% to 15% more than their different-sex counterparts.
The pay gap/premium is reduced quite significantly once the sample is limited to employed workers, which is unsurprising. As shown in Table O1 and Table O2, there are significant differences in labor supply for same-sex and different-sex workers. Men in male same-sex partnerships and women in different-sex partnerships are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
There is a significant decrease in the hourly wage gap now and an insignificant decrease in the annual wage gap. There is also a significant decrease in the annual earnings premium for women in female same-sex partnerships but no change in their hourly wage premium. This result is consistent with the labor supply results in Table O1 and O2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
O.6 Occupation Fixed Effects
We show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table O4 and O5.
In these tables, there is a smaller pay gap for men in male same-sex partnerships and a smaller pay premium for women in female same-sex partnerships when controlling for occupation or industry. However, qualitatively the results seem to be similar to the main results in Table O1 and Table O3. Men in male same-sex partnerships see a significant reduction in the labor supply gap, and women in female same-sex partnerships see a significant reduction in their labor supply and annual wage premium. There is no longer a significant effect on the wage gap for men in male same-sex partnerships. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
O.7 Sorting and Increased Reporting
One finding from Klawitter and Flatt [1998] suggests that LGB workers sort to areas with anti-discrimination laws. If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption. It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
Another potential concern would be increased reporting for same-sex partnerships. Same-sex partnerships, while less stigmatized in 2005-2019 than in previous years, were still heavily stigmatized. Anti-discrimination laws could signal to LGB people that their community is more accepting, and people may be more likely to declare that they are in a same-sex partnership. This change in reporting could affect the composition of our sample, which in turn could violate the identifying assumption.
To address this concern, we run the regression below to see if there is an increase in same-sex partnerships in a given county after an anti-discrimination law is passed:
$\text{SSP}_{ict}=\beta_{0}\text{Laws}_{ct}+\beta_{1}X_{i}+\beta_{2}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (6)
In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships. We present these results in Table O6.
We find no significant effect of anti-discrimination laws on the number of same-sex partnerships at the county level, suggesting there is minimal sorting or change in reporting following the passage of anti-discrimination laws.
Online Appendix Tables
Table O1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
Panel A: Men
(1) (2) (3) (4)
Labor Force Employed Labor Force Employed
Laws*SSP 0.013*** 0.014** 0.007 0.005
(0.005) (0.005) (0.005) (0.006)
SSP -0.074*** -0.079*** -0.068*** -0.073***
(0.003) (0.004) (0.003) (0.004)
State Laws Only X X
Observations 6,287,441 6,287,441 6,287,441 6,287,441
R-squared 0.135 0.114 0.135 0.114
Panel B: Women
(1) (2) (3) (4)
Labor Force Employed Labor Force Employed
Laws*SSP -0.013** -$0$016*** -0.013**
(0.006) (0.006) (0.006) (0.006)
SSP 0.081*** 0.073*** 0.080*** 0.072***
(0.004) (0.005) (0.004) (0.004)
State Laws Only X X
Observations 6,569,373 6,569,373 6,569,373 6,569,373
R-squared 0.091 0.088 0.091 0.088
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
Panel A: Men
(1) (2) (3) (4)
Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Laws*SSP 0.273* 0.089 0.180 -0.0744
(0.160) (0.080) (0.164) (0.0820)
SSP -2.196*** -0.575*** -2.106*** -0.477***
(0.114) (0.056) (0.0923) (0.0480)
State Laws Only X X
Observations 5,244,258 5,244,258 5,244,258 5,244,258
R-squared 0.034 0.011 0.034 0.011
Panel B: Women
(1) (2) (3) (4)
Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Laws*SSP -0.305** -0.184** -0.213 -0.171**
(0.136) (0.084) (0.134) (0.0790)
SSP 2.792*** 0.604*** 2.706*** 0.574***
(0.105) (0.066) (0.0924) (0.0577)
State Laws Only X X
Observations 4,366,603 4,366,603 4,366,603 4,366,603
R-squared 0.038 0.015 0.038 0.015
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O3: Effect of Anti-Discrimination Laws: Wages and Earnings
Panel A: Men
(1) (2) (3) (4) (5) (6) (7) (8)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP 0.028* 0.059 0.011 -0.041 0.0139 -0.00404 0.00639 -0.0581
(0.017) (0.050) (0.015) (0.035) (0.0184) (0.0525) (0.0168) (0.0373)
SSP -0.244*** -0.757*** -0.082*** -0.114*** -0.233*** -0.716*** -$0**$112***
(0.012) (0.038) (0.009) (0.023) (0.0108) (0.0354) (0.00745) (0.0184)
Employed Only X X X X
State Laws Only X X X X
Observations 6,287,441 6,287,441 5,244,258 5,244,258 5,244,258 6,287,441 5,244,258 6,287,441
R-squared 0.157 0.121 0.127 0.041 0.127 0.157 0.041 0.121
Panel B: Women
(1) (2) (3) (4) (5) (6) (7) (8)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP -0.019 -0.129** -0.024** -0.062** -0.0203 -0.153** -0.0144 -0.0453
(0.020) (0.066) (0.011) (0.029) (0.0202) (0.0668) (0.0112) (0.0297)
SSP 0.220*** 0.815*** 0.053*** 0.149*** 0.218*** 0.809*** 0.0448*** 0.132***
(0.014) (0.049) (0.007) (0.021) (0.0116) (0.0409) (0.00635) (0.0180)
Employed Only X X X X
State Laws Only X X X X
Observations 6,569,373 6,569,373 4,366,603 4,366,603 6,569,373 6,569,373 4,366,603 4,366,603
R-squared 0.136 0.106 0.123 0.036 0.136 0.106 0.123 0.036
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. Columns (3) and (4) are limited to those who are employed. All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Columns (5) - (8) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
Men Women
(1) (2) (1) (2)
Labor Force Employed Labor Force Employed
Laws*SSP 0.009*** 0.009** -0.009*** -0.009***
(0.003) (0.004) (0.003) (0.004)
SSP -0.037*** -0.045*** 0.039*** 0.036***
(0.002) (0.003) (0.002) (0.003)
Observations 6,287,441 6,287,441 6,569,373 6,569,373
R-squared 0.372 0.461 0.517 0.461
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients show the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients give the labor supply gap or premium. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O5: Anti-Discrimination Laws on Pay with Occupation FEs
Panel A: Men
(1) (2) (3) (4)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP 0.013 0.018 0.011 -0.028
(0.013) (0.035) (0.013) (0.033)
SSP -0.113*** -0.351*** -0.043*** -0.051**
(0.008) (0.025) (0.008) (0.022)
Employed only X X
Observations 6,287,441 6,287,441 5,244,258 5,244,258
R-squared 0.372 0.372 0.248 0.169
Panel B: Women
(1) (2) (3) (4)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP -0.015 -0.104*** -0.024*** -0.063**
(0.011) (0.034) (0.009) (0.025)
SSP 0.117*** 0.437*** 0.056*** 0.172***
(0.008) (0.028) (0.007) (0.018)
Employed only X X
Observations 6,569,373 6,569,373 4,366,603 4,366,603
R-squared 0.513 0.527 0.303 0.233
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O6: Sorting and Reporting
(1) (2) (3) (4) (5)
All SSP Male SSP Female SSP Age Years of Education
ENDA -0.0002 0.0004 -0.0008 0.0257 -0.0212
(0.0005) (0.0005) (0.0005) (0.0648) (0.0182)
Observations 16,071,420 7,850,821 8,220,599 16,071,420 16,071,420
Mean of DV 0.0122 0.0123 0.0122 45.5220 13.8539
$R^{2}$ 0.0043 0.0081 0.0024 0.0101 0.0467
  • Notes: Data comes from the 2005-2019 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law. The first column looks at both men and women with the next two columns separating the sexes. Column (4) investigates age and (5) looks at years of education. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O7: Differential Responses between Male and Female Same-Sex Partnerships
(1) (2) (3) (4)
One Earner HH Diff in Hours Worked Any Children Number of Children
Laws*FemSSP -0.00308 0.947*** 0.0344*** 0.0544***
(0.00868) (0.364) (0.00846) (0.0195)
FemSSP 0.00445 -0.729** 0.123*** 0.193***
(0.00717) (0.317) (0.00743) (0.0176)
Observations 73,181 73,181 73,181 73,181
R-squared 0.046 0.043 0.122 0.111
  • Notes: Data comes from the 2005-2016 yearly ACS comparing women in same-sex partnerships to men in same-sex partnerships. The first column examines if the household has only one-earner. The second column tests the difference in absolute terms of hours worked between the two partners. The third and fourth columns examines how households differ with having children. The first row of coefficients show the differential effect of anti-discrimination between lesbian and gay households. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.